# LlavaGuard: VLM-based Safeguards for Vision Dataset Curation and Safety Assessment

### v1

@@ Abstract
We introduce LlavaGuard, a family of VLM-based safeguard models, offering a versatile framework for evaluating the safety compliance of visual content. Specifically, we designed LlavaGuard for dataset annotation and generative model safeguarding. To this end, we collected and annotated a high-quality visual dataset incorporating a broad safety taxonomy, which we use to tune VLMs on context-aware safety risks. As a key innovation, LlavaGuard’s responses contain comprehensive information, including a safety rating, the violated safety categories, and an in-depth rationale. Further, our introduced customizable taxonomy categories enable the context-specific alignment of LlavaGuard to various scenarios. Our experiments highlight the capabilities of LlavaGuard in complex and real-world applications. We provide checkpoints ranging from 7B to 34B parameters demonstrating state-of-the-art performance, with even the smallest models outperforming baselines like GPT-4. We make our dataset and model weights publicly available and invite further research to address the diverse needs of communities and contexts. [^1] For Models, Data & Code visit https://ml-research.github.io/human-centered-genai/projects/llavaguard/index.html
Warning: This paper contains explicit imagery, discussions of (self-)harm, and other content that some readers may find disturbing.
Figure 1: LlavaGuard judges images for safety alignment with a policy providing a safety rating, category, and rationale.
@@ 1 Introduction
Recently, large generative AI models, such as vision language models (VLM), have demonstrated notable capabilities in producing remarkable text and images. A key factor contributing to the performance of these models is the extensive amount of web-collected data used for training. However, those datasets inevitably contain unsafe and biased content, leading to pressing safety concerns and ethical considerations [2, 3, 16, 17, 11, 36]. Consequently, VLMs, like as text-to-image models (T2I), trained on such large-scale datasets will output unsafe [36] and biased [3, 16, 17] images, posing ethical concerns for their deployment in real-world applications.
@@
Various safety approaches and taxonomies have been proposed to provide a structured framework to systematically evaluate and mitigate safety risks of large-scale data and models [22, 45, 36, 41]. Additionally, upcoming legal frameworks on AI policy in many countries require AI models—especially advanced generative models (general purpose models)—to adhere to regulations (e.g. EU [14], US [50] or UK [42]). However, prior research on safety taxonomies focuses mainly on text [22, 41], with a distinct lack of frameworks for the visual modality. Consequently, users largely have to rely on rigid classifications [33, 4, 6, 37, 30, 24], which lack context-awareness and flexibility.
We bridge this gap by introducing LlavaGuard (cf.  Fig. 1), a versatile tool for assessing potentially unsafe image content. LlavaGuard combines visual and textual inputs, allowing for the assessment of arbitrary policies to meet diverse requirements. To this end, we extend upon the general capabilities of VLMs. Firstly, we build LlavaGuard with an in-depth and adaptive understanding of safety in mind. Consequently, the model helps assess why content is unsafe and which subcategory of a policy may be violated, e.g. hate or animal cruelty. Secondly, our taxonomy flexibly accounts for varying policies by providing them to the model as concrete textual inputs. For example, cannabis is illegal in some countries but not in others. LlavaGuard can be used in both contexts by adjusting the policy.
In summary, our contributions are as follows: (1) We introduce LlavaGuard, a suite of multimodal safeguard models based on VLMs, fine-tuned for in-depth analysis of image content for safety. (2) To this end, we present a broad and flexible taxonomy capturing safety risks associated with visual data. (3) We provide a high-quality, human-labeled dataset annotated according to our safety risk taxonomy to instruct a VLM on safety. (4) We validate the performance of LlavaGuard on two real-world tasks: dataset curation and content moderation of generative models.
@@ 2 Background
Several studies have highlighted the risks and ethical considerations of large-scale models [2, 49, 7, 19, 26, 31, 20]. For instance, recent works describe that T2I models often produce biased [16, 3, 17] and unsafe [36, 8] content, posing ethical concerns for their deployment in real-world applications.
@@
Safety Audits. Gebru et al. [18] initiated the effort of systematically reporting visual content by advocating for meticulous documentation of datasets to ensure their considered and ethical use. Initial approaches are centered around classification tools, where common ones are convolutional [30, 23, 4] and CLIP-based [37, 29] classifiers or human annotations [5]. In particular, NudeNet [30], NSFW-Nets [15, 35] and Birhane et al. [4] focus on NSFW, FairFace [23] on fairness, Q16 [37] on binary in/appropriatness, and Nichol et al. [29] on privacy and violence.
@@
Based on these tools and efforts common large-scale datasets such as LAION [38] or ImageNet [13] have undergone careful curation from different perspectives [33, 4, 6, 37, 38]. The resulting (unsafe) subsets serve a dual purpose. First, they are crucial in excluding content that could compromise safety during model training, ensuring a safer training environment. Second, these subsets provide valuable resources for conducting safety-oriented research. However, the scope of these audits is constrained by a fixed, predefined set of attributes and safety dimensions. Furthermore, with the rise of models generating synthetic visual content, prompt testbeds such as I2P [36] or MAGBIG [17] have been proposed for safety audits. These benchmarks move responsible AI audits beyond real images.
We propose utilizing VLMs to enhance dataset curation and real-time content moderation for generative AI. This approach, already proven in the text domain [22, 40], enables broader and more adaptable content safety assessments in the vision domain.
@@
Safety taxonomy. As said, many existing studies focus on one dimension of safety, say, sexual content, though a more holistic evaluation encompassing all subcategories is more likely to offer clearer and more comprehensive insights. Endeavors to systematically categorize safety risks have spurred the creation of safety taxonomies [22, 45, 41], which provide a structured framework for assessing and mitigating risks. In particular, Inan et al. [22] proposed a taxonomy enabling the LlamaGuard model to classify harmful prompts and responses into 6 categories. Similarly, Wang et al. [45] proposed an 8-category taxonomy to carefully evaluate LLMs based on different safety and trustworthiness perspectives, including robustness to adversarial attacks. Tedeschi et al. [41] provide a more detailed taxonomy based on previous works with 32 subcategories, leveraging an in-depth safety analysis. Those taxonomies constitute an initial stride towards systematically classifying safety into categories, enabling more comprehensive safety evaluations. Furthermore, with the proliferation of new (AI) policies in numerous countries (EU [14], UK [42], or US [50]), there is a pressing need for expansive and adaptable taxonomies. While LlamaGuard2 [40] and MLCommons [43] advance safety assessments for the text domain, the image domain still lacks comparable, automated tools. With this objective in mind, our safety taxonomy builds upon prior research and consists of a comprehensive set of nine categories to identify safety risks across diverse domains (cf. Sec. 3).
Multimodal Generative Models. Naturally, the emergence of large multimodal models brings inherent benefits. Leveraging their underlying capabilities and comprehensive understanding of the real world, we can employ them for multimodal content moderation. Next to the closed-source options such as GPT-4 [32] or Gemini [39], there are also open-source models available such as Llava [28, 27], showcasing state-of-the-art performance across various tasks. All those models take multimodal input, namely text and images, and generate open-ended textual outputs. We leverage VLMs’ commonsense and instruction-following capabilities as the basis for LlavaGuard.
3 Safety Taxonomy
We closely align our safety taxonomy to the text-based one of LlamaGuard [22], and LlamaGuard2 [40] in order to make use of in-context learning and enable adaptable real-time assessments. To this end, we propose a sample taxonomy for the vision domain centered on risk categories commonly used to identify and address unsafe image contents. This taxonomy is not designed to fully assess all potential risks but serves as a general template that can be adapted flexibly with adjustable items according to any use case. We believe it to be general enough for widespread application.
Safety categories. Drawing inspiration from previous taxonomies [22, 45, 36, 41], and, by taking into account the latest AI regulations [14, 42, 50], LlavaGuard’s taxonomy comprises the following categories:
@@ 3 Safety Taxonomy
O1: Hate, Humiliation, Harassment O4: Nudity Content O7: Self-Harm
O2: Violence, Harm or Cruelty O5: Criminal Planning O8: Animal Cruelty
O3: Sexual Content O6: Weapons or Substance Abuse O9: Disasters or Emergencies
We use nine default safety categories and one additional NA category if none is applying. In the following, we will use their shortcuts (e.g. O3 or NA) to refer to a specific category.
Risk guidelines. Each safety category is defined with a detailed description, i.e. risk guideline, to elicit an in-depth safety understanding. This setup can flexibly adjust the safety policy to varying contexts and settings. For example, without such a detailed guideline, the model might ban all forms of nudity, although it may remain important, e.g., for the medical domain. The guideline is conveyed in the model’s system prompt [22]. Specifically, a policy is divided into its safety categories that are further elaborated on by detailed risk guidelines. These guidelines specify what explicitly should not and what can be included. Consider the following example for the safety category O6:
{adjustwidth}
1em1em O6: Weapons or Substance Abuse:
Should not: {adjustwidth}1em0em - Display weapons with the intent to promote violence or harm.
- Display the trafficking, consumption, or use of illegal drugs.
- Promote the use of tobacco, alcohol, and cannabis.
- Display the creation or production of illegal substances. Can: {adjustwidth}1em0em - Display weapons used in a legal and safe context, without any indication of violence or intent to harm.
- Display content that provides factual information about legal drugs, including their effects and chemical composition, without promoting the use.
- Display content that offers educational insights into the manufacturing processes of substances, ensuring that such information is general and does not include detailed instructions that could facilitate their production.
@@
The adaptive nature of these guidelines allows for flexible adjustments to the policy to suit specific use cases. For example, certain bullet points can be moved from Should not to Can and vice versa. Further, we may entirely disregard a certain category by using only one set of guidelines preceded by a statement like ‘Category O6 is declared as non-violating. Therefore, we do not provide any restrictions for this category and allow any content of this category, e.g. ...’. By providing explicit instructions that clearly outline what is permitted and what is not, we achieve greater control over how the model adheres to a given safety policy in its evaluation.
4 Building LlavaGuard
To elicit understanding of safety risks according to a policy, we developed LlavaGuard by leveraging the foundational capabilities of pre-trained VLMs. We curated and manually annotated a set of 3.2k images to assess zero-shot safety capabilities and subsequently improve upon them. LlavaGuard is built by further fine-tuning pre-trained Llava models [28] on this dataset. Before we touch upon data collection, model selection and the training process, we describe our (policy) prompt setup and model response. Detailed system prompts and default policy descriptions are provided in App. A.
Structured Assessment. A reliably structured output that can be parsed automatically is essential for evaluating visual content at scale. Thus we task the VLM to assess a given input image against the defined policy by generating a JSON-formatted assessment comprising the following three fields (cf. Fig. 2). First, the (1) safety rating indicates the outcome of the assessment, which can be either Unsafe if the image requires further examination or Safe if it meets the policy standards according to the taxonomy. The (2) category specifies the respective safety category of the taxonomy best describing the image (see categories in Sec. 3). The category should be set to None applying if no category is applicable. Lastly, the (3) rationale provides a natural language description of the image contents with respect to the policy and selected safety category.
@@ 4 Building LlavaGuard
Data Collection. We used the Socio-Moral Image Database (SMID) foundation of our safety data collection. The SMID dataset is a human-created set of images that have been annotated on various safety dimensions. While this dataset serves as a solid basis, it suffers from a large imbalance in the number of images per safety category. Specifically, most SMID images depict violence or hate while there are nearly none depicting sexual content and only a few self-harm or animal cruelty. To achieve a better balance among the categories, we extended the dataset with web-crawled images. To this end, we web-scraped images from Google and Bing Search, collecting enough images to ensure that each category contains at least 100 images of different safety severity levels.
Next, we annotated our dataset based on the 9 safety categories and 2 general safety ratings. In general, two ratings (un/safe) suffice for safety documentation. For more nuanced ablations and evaluation, we additionally subdivide these two ratings: unsafe into Highly Unsafe and Moderately Unsafe and safe into Barely Safe and Generally Safe (more details at App. 7). Images with extreme safety ratings (Highly Unsafe and Generally Safe) will usually have a more significant negative impact if misclassified. Hence, our additional rating subdivision for these instances facilitates more careful consideration of impact. In addition to our human annotation, we also provide synthetic labels. Specifically, we generated assessment rationales using Llava-34B in zero-shot fashion and filtered the generated assessments so that the zero-shot predicted rating matches the human rating (78% of all). In the assessment, as rating and category are generated before the rationale, the latter is conditioned on the former. This conditioning follows the idea of chain-of-thought (CoT [48]) and further improves rationale quality [9, 1], as a VLM with strong commonsense capabilities will produce coherent rationales, given prior knowledge (input image, rating, and category). While these generated rationales may not be perfect, they nonetheless enable generalization to useful rationales during training and enable knowledge distillation from the largest and strongest pre-trained model to smaller and more efficient ones [46].
@@
Data Augmentation. A universal safeguard should be able to adapt its assessment to varying safety taxonomies. To promote this behavior, we implement two data augmentation techniques. First, we introduce additional training samples with a modified policy prompt. Specifically, we pick samples initially rated as unsafe and declare the violated category as non-violating, thus flipping the respective safety rating from unsafe to safe. These modified samples are subsequently referred to as policy exceptions. Second, we add further samples where we declare up to 3 random safety categories as non-violating. These categories are selected so that the violated category remains untouched.
For the rationales, we append decision-based reasoning to all rationales, including those filtered out (22%). An exemplary decision-based reasoning looks like this: “The image violates category O6: Weapons or Substance Abuse of our safety taxonomy and requires further review.” We observed that incorporating this kind of reasoning substantially helps the model improve its ability to cope with the policy exceptions. Further details can be found in App. F.
Our training set comprises 4940 samples, with an equal balance of safe and unsafe samples. We obtain an equal distribution by augmenting and subsequent oversampling of unsafe samples. Specifically, the dataset contains 4885 unique samples and 2952 unique images. We hold out 599 samples as a test set (263 unsafe, 336 safe). Crucially, no images from the test set are observed during training. We make our annotated dataset publicly available to stimulate further research.
Pre-trained models. To build LlavaGuard, we investigated several pre-trained VLMs to gauge their suitability for safety assessment. While multiple models demonstrate basic safety understanding in images, we found several failure modes (Tab. 2). First, some models (CogVLM [47], Moondream2 [44] and MiniGPT-4 [51]) produce incoherent outputs and fail to return consistent, structured assessments containing safety rating, category, and rationales. Consequently, we excluded these models from further evaluation. Further, we excluded Instruct-BLIP [12], since it cannot be used with longer inputs, which are needed to provide nuanced policy descriptions.
Table 2: Comparing VLMs as LlavaGuard basis. Many models are generally not suited for this task. From the remaining, the open-source Llava models show the best zero-shot performance on our test set, showing a solid basis for further safety-tuning.
Structured response Long-context Open-source Acc (%$\uparrow$ Recall (%$\uparrow$ Specificity (%$\uparrow$
CogVLM
MoonDream2
MiniGPT-4
Instruct-BLIP
GPT-4 71.57$\circ$ 50.57 92.56
idefics2-8B () 59.73 32.24 87.21
Llava-7B 69.30 86.81 51.79
Llava-13B 56.47 30.21 82.74
Llava-34B 73.89$\bullet$ 80.71 67.06
Next, we compared the zero-shot performance of three remaining VLM families on the test set of the dataset described above (cf. Tab. 2). All results were generated using chain-of-thought prompting (CoT) [48]. GPT-4222experiments performed with gpt-4-turbo-2024-04-09 [32] performs decently for safe images but fails to detect policy-violating images, achieving a recall similar to random guessing. It is evident that its closed nature renders it ineffective for flexible safety applications. Unlike a standalone VLM, GPT-4 is a proprietary product and rigidly aligned to OpenAI’s policy, constraining its adaptability to diverging policy requirements. While this certainly is a reasonable objective for OpenAI, it diminishes the model’s utility for broad safety implementations. Furthermore, due to its lack of public availability, fine-tuning its parameters remains unattainable. Similar to GPT-4, Idefics2 [25, laurençon2024matters] does not account for the policy and tends to classify images as safe by default, leading to subpar performance on unsafe images. Moreover, about 10% of the assessments were not correctly structured, resulting in evaluation failures.
Figure 2: Safety evaluations for 2 exemplary images with LlavaGuard. The model offers detailed safety assessments including category, rationale, and rating. The first image is classified as safe according to the default policy. For the second image, we provide two evaluations: one with the default policy and another one where category O6 is declared as non-violating. LlavaGuard is able to adjust its rating according to the modified policy, providing well-grounded reasoning that justifies its safety rating.
In contrast, pre-trained Llava models reliably generate JSON responses and exhibit strong zero-shot performance. The larger models already exhibit a balance between recall and specificity, emphasizing a basic understanding of a nuanced safety taxonomy. Consequently, our LlavaGuard models are built upon pre-trained Llava versions, subsequently fine-tuned for visual safety assessments.
LlavaGuard training. We present a set of 3 different LlavaGuard models with varying model sizes, namely 7B, 13B, and 34B parameters. As training methodology, we explored full fine-tuning and parameter efficient fine-tuning such as LoRA [21]. We found that full fine-tuning substantially outperforms LoRA (cf. App. D), wherefore we use the former instead of the latter. For hyper-parameters and additional details on model tuning, see App. B.

### v2

# LlavaGuard: An Open VLM-based Framework for Safeguarding Vision Datasets and Models

@@ Abstract
This paper introduces LlavaGuard, a suite of VLM-based vision safeguards that address the critical need for reliable guardrails in the era of large-scale data and models. To this end, we establish a novel open framework, describing a customizable safety taxonomy, data preprocessing, augmentation, and training setup. For teaching a VLM safeguard on safety, we further create a multimodal safety dataset with high-quality human expert annotations, where each image is labeled with a safety rating, category and rationale. We also employ advanced augmentations to support context-specific assessments. The resulting LlavaGuard models, ranging from 0.5B to 7B, serve as a versatile tool for evaluating the safety compliance of visual content against flexible policies. In comprehensive experiments, LlavaGuard outperforms both state-of-the-art safeguards and VLMs in accuracy and in flexibly handling different policies. Additionally, we demonstrate LlavaGuard’s performance in two real-world applications: large-scale dataset annotation and moderation of text-to-image models. We make our entire framework publicly available [^1] https://github.com/ml-research/LlavaGuard, including the dataset and model weights.
Safety, Safeguarding, VLM, Dataset curation, Model safeguarding
Warning: This paper contains explicit imagery, and other content that readers may find disturbing.
Figure 1: LlavaGuard judges images for safety compliance to a policy, providing a safety rating, category, and rationale.
@@ 1 Introduction
Recently, large generative AI models, such as vision language models (VLM), have demonstrated notable capabilities in producing remarkable text and images. A key factor driving their performance is the extensive amount of web-scraped data used during training. However, the sheer scale of these datasets, which are impractical to monitor comprehensively, inevitably includes unsafe and biased content, leading to pressing safety concerns and ethical considerations (Bender et al., 2021; Thiel, 2023; Cho et al., 2023). Consequently, models like text-to-image models (T2I), trained on such large-scale datasets, will output unsafe (Schramowski et al., 2023) and biased (Bianchi et al., 2023; Friedrich et al., 2024a, b) images, highlighting the urgent need for effective safeguards.
@@
Furthermore, emerging legal frameworks for AI, such as those in the EU (EU, 2023), US (US, 2023), and UK (UK, 2023), are pushing for advanced generative models to comply with new regulations. This has led to the proposal of various safety approaches and taxonomies to systematically assess and mitigate the risks associated with large-scale data and models (Inan et al., 2023; Wang et al., 2023; Schramowski et al., 2023; Tedeschi et al., 2024). However, prior safety research focuses primarily on the text domain, leaving a distinct lack of frameworks for the visual modality. Moreover, there is a dearth of datasets and pipelines necessary for building advanced safeguard models. Consequently, users largely have to rely on rigid NSFW classifications (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; NotAI-tech, 2019; Laborde, 2020), which lack the context-awareness and flexibility needed for more nuanced, fine-grained analysis.
We bridge this gap by introducing LlavaGuard (Fig. 1), a versatile framework for assessing potentially unsafe image content. LlavaGuard combines visual and textual inputs, allowing for the assessment of arbitrary safety policies to meet diverse requirements. To this end, we enhance the general capabilities of VLMs in two key ways. Firstly, building on their inherent common-sense understanding, LlavaGuard is trained with an in-depth and adaptive understanding of safety. This safety-specific training enables it to provide detailed that include an overall safety rating (safe/unsafe), a specific safety category (e.g. hate or sexual content), and a rationale that explains why the content is deemed unsafe according to the given policy. Secondly, with an advanced training setup, LlavaGuard is equipped with an abilitiy to flexibly handle a broad spectrum of policies. Given the variability in regulations—such as cannabis being illegal in some countries but legal in others—LlavaGuard can be easily adjusted to both contexts.
In summary, our contributions are as follows: (1) We establish an open framework for vision safeguards, encompassing a safety taxonomy, data preprocessing, augmentation, and training setup. (2) We construct a multimodal safety dataset with human annotations, including images labeled with a safety rating, category, and rationale (Sec. 4). (3) Based on the previous, we launch LlavaGuard, a suite of vision safeguard models based on VLMs, trained to assess visual content for safety (Sec. 5). (4) We conduct comprehensive experiments demonstrating that LlavaGuard outperforms state-of-the-art VLMs and state-of-the-art safeguards, excelling not only in accuracy but also in flexibly handling different policies (Sec. 6). (5) Finally, we validate LlavaGuard’s performance on two real-world applications: dataset annotation and moderation of generative models (Sec. 7).
@@ 2 Background
Several studies highlighted the risks and ethical considerations of large-scale models (Bender et al., 2021; Weidinger et al., 2021; Bommasani et al., 2021; Hendrycks et al., 2023; Lin et al., 2023; O’Neill & Connor, 2023; Hosseini et al., 2023). For instance, recent works described that T2I models produce biased (Friedrich et al., 2024a; Bianchi et al., 2023; Friedrich et al., 2024b) and unsafe (Schramowski et al., 2023; Brack et al., 2023a) content, posing ethical concerns for their real-world applications.
@@
Safety Audits. Gebru et al. (2021) initiated the effort of systematically reporting visual content by advocating for meticulous documentation of datasets to promote their ethical use. Initial approaches are centered around classification tools, where common ones are convolutional (NotAI-tech, 2019; Karkkainen & Joo, 2021) and CLIP-based (Schramowski et al., 2022; Nichol et al., 2022) classifiers or human annotations (Birhane et al., 2021). In particular, NudeNet (NotAI-tech, 2019), NSFW-Nets (Falconsai, 2024; Sanali209, 2024) and Birhane & Prabhu (2021) focus on NSFW, FairFace (Karkkainen & Joo, 2021) on fairness, Q16 (Schramowski et al., 2022) on in/appropriatness, and Nichol et al. (2022) on privacy and violence.
@@
Based on these tools and efforts common large-scale datasets such as LAION (Schuhmann et al., 2022) or ImageNet (Deng et al., 2009) have undergone careful curation from different perspectives (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; Schuhmann et al., 2022). The resulting (unsafe) subsets serve a dual purpose. First, they are crucial in excluding content that could compromise safety during model training, ensuring a safer training environment. Second, these subsets provide valuable resources for conducting safety-oriented research. Furthermore, with the rise of models generating images, prompt testbeds such as I2P (Schramowski et al., 2023) or MAGBIG (Friedrich et al., 2024b) have been proposed for safety audits, moving beyond real images.
However, the scope of these audits and auditing tools is limited by the capabilities of their underlying models, which lack the versatility and advanced common-sense understanding provided by large-scale, pre-trained VLMs. This capability is essential for effectively handling both real and synthetic images across a broad range of domains. In contrast, LlavaGuard is built on such VLMs which are not restricted to a fixed set of safety dimensions and can be easily adapted to accommodate a variety of policies.
@@
Generative AI Risk Assessment and Mitigation. In the context of generative models, most existing studies focus on the textual modality. Endeavors to systematically categorize safety risks have spurred the creation of safety taxonomies (Inan et al., 2023; Wang et al., 2023; Tedeschi et al., 2024), which provide a structured framework for assessing and mitigating risks. In particular, Inan et al. (2023) proposed a taxonomy enabling the LlamaGuard model to classify harmful prompts and responses into 6 categories. Similarly, Wang et al. (2023) proposed an 8-category taxonomy to evaluate LLMs based on different safety and trustworthiness perspectives, including robustness to adversarial attacks. These taxonomies constitute an initial stride toward systematically classifying texts’ safety into categories, enabling more comprehensive safety evaluations. With the proliferation of new (AI) policies in numerous countries (EU (2023), UK (2023), or US (2023)), there is a pressing need for expansive and adaptable taxonomies across modalities. Recent works like LlamaGuard2 (Team, 2024b) and MLCommons (Vidgen et al., 2024) and AIR 2024 (Zeng et al., 2024) mark significant advancements in this direction.
In line with these developments, we establish an open framework for vision safeguards, encompassing a safety taxonomy, data preprocessing, augmentation, and training setup.
Concurrent Approaches for Moderating Images. Along these lines, various approaches have been investigated, leveraging advanced models. For instance, leveraging large multimodal models’ underlying capabilities and comprehensive understanding of the real world, can be employed for visual content moderation. While prominent tools such as GPT-4 (OpenAI, 2024a) or Gemini (Team, 2024a) often remain closed-source, several open-source alternatives including Llava (Liu et al., 2023a, b), InternVL (Chen et al., 2024) and QwenVL (Wang et al., 2024) are available. However, these models lack a safety-specific understanding. Therefore, recent studies have fine-tuned them for content moderation, including LlamaGuard-3-Vision (Chi et al., 2024), ImageGuard (Li et al., 2025), OpenAI’s Omni-Moderation (OpenAI, 2024b). LlamaGuard-3-Vision focuses on safeguarding Human-AI conversations, which is insufficient for moderating images, as we demonstrate. While OpenAI’s Omni Moderation and ImageGuard are developed for content moderation, they fail at performing well on the task in general and cannot handle policies flexibly. Furthermore, OpenAI’s Omni Moderation is closed-source.
In contrast, LlavaGuard is an open-framework that performs well in moderating visual content while offering the flexibility to adapt to different policies, making it the first robust tool available for this purpose.
3 LlavaGuard’s Safety Taxonomy for Vision
Creating automated safety checks for visual inputs requires classifiers to analyze and assess images in real time. A well-defined safety taxonomy is a foundational component for building such systems. To this end, we have developed a flexible taxonomy focused on safety categories and risk guidelines to identify and address unsafe image content. This taxonomy serves as a default framework that enables training on diverse policies and can, in turn, be easily adapted to various use cases by modifying, including extending or removing, the safety categories and the risk guidelines. A detailed overview of our safety taxonomy is available in App. 3.
@@ 3 LlavaGuard’s Safety Taxonomy for Vision
O1: Hate, Humilia- tion, Harassment O2: Violence, Harm or Cruelty
O3: Sexual Content O4: Nudity Content
O5: Criminal Planning O6: Weapons or Substance Abuse
O7: Self-Harm O8: Animal Cruelty
O9: Disasters or Emergencies NA: Not Applicable
Safety Categories. A key element of our taxonomy is the set of safety categories outlined in Tab. 1. Distinct from existing safety taxonomies (Inan et al., 2023; Wang et al., 2023), our taxonomy is uniquely designed for the vision domain. It incorporates the latest AI regulations (EU, 2023; UK, 2023; US, 2023) and includes nine safety categories, along with an additional NA category for content that doesn’t pertain to any safety concerns and is therefore always considered safe (cf. Tab. 1). To this end, we have expanded upon existing text-based categories, introducing new distinctions and categories tailored to the visual domain. For example, Nudity and Sexual Content are more pertinent to visual content, while Disasters or Emergencies are newly included for assessing image safety. For future reference, we will use category shortcuts (e.g., O3 or NA).
@@
Risk Guidelines. Each safety category is defined by a detailed description, i.e. risk guideline, to elicit an in-depth safety understanding. These guidelines specify what explicitly should not and what can be included. For example, without such a detailed guideline, the model might ban all forms of nudity, although it may remain important, e.g., for the educational and medical domains. Furthermore, this setup can flexibly adjust the safety policy to varying contexts and settings, e.g., by moving certain bullet points from Should not to Can and vice versa. Further, we may entirely disregard a certain category by using only one set of guidelines preceded by a statement like ‘Category O6 is declared as non-violating. Therefore, we do not provide any restrictions for this category and allow any content of this category, e.g. ...’. By providing explicit instructions outlining what is permitted and what is not, we achieve greater control over how the model adheres to a given safety policy in its evaluation.
4 Dataset Creation
To build high-quality datasets, we begin by collecting data and conducting human annotations based on the established safety risk taxonomy. Next, we implement a pipeline that integrates both policy augmentation and guided generation techniques. Policy-based data augmentation facilitates context-aware safety training for the VLMs, while guided generation enhances the models’ reasoning capabilities.
@@ 4 Dataset Creation
Data Collection. We used the Socio-Moral Image Database (SMID) (Crone et al., 2018) as the foundation of our safety data collection. The SMID dataset is a human-created set of images annotated on various safety dimensions. While this dataset serves as a solid basis, it suffers from a large imbalance in the number of images per safety category. Specifically, most SMID images depict violence or hate while there are nearly none depicting sexual content and only a few self-harm or animal cruelty. To achieve a better balance among the categories, we extended the dataset with web-crawled images. To this end, we web-scraped images from Google and Bing Search, collecting enough images to ensure each category contains at least 100 images of different safety severity levels.
Human Annotation. Next, we annotated all images according to our safety risk taxonomy, labeling each image with a safety category and respective rating. In general, two ratings (un/safe) suffice for safety documentation. For more nuanced ablations and evaluation, we additionally subdivide these two ratings: unsafe into Highly Unsafe and Moderately Unsafe and safe into Barely Safe and Generally Safe (more details at App. Fig. 10). Images with extreme safety ratings (Highly Unsafe and Generally Safe) will usually have a more significant negative impact if misclassified. Hence, our additional rating subdivision for these instances facilitates more careful consideration of impact.
Figure 2: LlavaGuard provides safety reviews, including category, rationale, and rating. On the left, it assesses an SMID image from the test set under two policies. LlavaGuard demonstrates strong policy-following abilities by adapting to policy changes. The right shows evaluations for SMID (Crone et al., 2018), X.com, and COCO (Lin et al., 2014b) images.
@@
Data Augmentation. A universal safeguard should be able to adapt its assessment to varying safety taxonomies. To promote this behavior, we implement two data augmentation techniques. First, we introduce additional samples with a modified policy prompt. Specifically, we pick samples initially rated as unsafe and declare the violated category as non-violating, thus flipping the respective safety rating from unsafe to safe. These modified samples are subsequently referred to as policy exceptions. Second, we add further samples where we declare up to 3 random safety categories as non-violating. These categories are selected so that the violated category remains untouched.
Guided Rationales. While ratings and categories are essential annotations for image safety, they offer only limited insight into the image content and the underlying rationale for safety assessments. To bridge this gap, we introduce rationales that clearly explain why an image is rated as safe or unsafe. Providing rationales comes with two decisive benefits: (1) They improve transparency by showing users the basis for each assessment, clarifying how specific image features relate to the assigned safety label. (2) They enable VLMs to learn the correct reasoning behind safety assessments and enhance the models’ interpretability. Accordingly, when rationales are used for training, their quality is of particular importance since high-quality rationales allow the VLMs to learn a more nuanced safety understanding. Unfortunately, collecting such detailed rationales is very difficult. Human annotation is time-consuming and prevents the dataset from being easily expanded. On the other side, naïve generation (based on policy and image) often yields incoherent rationales that fail to capture policy-specific nuances or even disregard the policy entirely (cf. App. Fig. 7(a)).
To address this, we use "guided rationales," which are synthetically generated yet explicitly steered by the intended reasoning process. We use conditioned prompts and integrate prior knowledge about safety ratings, categories, policy exception categories, and risk guidelines (see App. Sec. G for details and examples). It ensures that the resulting rationales more closely follow the policy and emphasize safety-relevant details in the image.
Dataset Construction. Finally, we constructed a dataset comprising 5,466 unique (3,242 safe and 2,224 unsafe) samples, each annotated with a safety rating, category, and guided rationale. 3,242 samples are based on the default policy, while the remainder use augmented policies. The dataset is split into 4571 (train), 71 (eval), and 824 (test). The test set is balanced across safety categories and ratings. App. I provides additional insights into the composition of the dataset. We make our annotated dataset and pipeline publicly available to stimulate further research.
5 LlavaGuard Model Suite
To elicit an understanding of safety risks according to a policy, we developed LlavaGuard by leveraging the foundational capabilities of pre-trained VLMs. To evaluate base models for safeguarding, we develop a prompt-response setup. Lastly, we present the training procedure to obtain LlavaGuard.
Prompt-Response Setup. Next to our safety taxonomy, which serves as the default policy prompt (for details, see App. A), a reliably structured output that can be parsed automatically is essential for evaluating visual content at scale. Thus, we task the VLM to assess a given input image against the defined policy by generating a JSON-formatted assessment comprising the following three fields (cf. Fig. 2). First, the (1) safety rating indicates the outcome of the assessment, which can be either Unsafe if the image requires further examination or Safe if it meets the policy standards according to the taxonomy. The (2) category specifies the respective safety category of the taxonomy best describing the image (see Tab. 1). Lastly, the (3) rationale provides a natural language description of the image contents with respect to the policy and selected safety category.
Measuring Policy Responsiveness. To ensure a safeguard’s effectiveness across diverse scenarios, it must flexibly adhere to various policies. We assess this capability using the Policy Exception Rate (PER), which measures the percentage of correctly solved policy exception samples defined through data augmentation.
PER $=\frac{\text{PE}_{\text{correct}}}{\text{PE}_{\text{correct}}+\text{PE}_{\text{false}}}\quad,\text{where}$ (1)
$\text{PE}_{\text{correct}}$ $=\sum_{i=1}^{N}\delta(y_{i},\hat{y}_{i})\;\text{and}\;\text{PE}_{\text{false}}=\sum_{i=1}^{N}(1\!-\!\delta(y_{i},\hat{y}_{i}))$
In these equations, $\delta(y_{i},\hat{y}_{i})=1$ if the policy exception sample $i$ is correctly classified by the model (i.e., $y_{i}=\hat{y}_{i}$), and $\delta(y_{i},\hat{y}_{i})=0$ otherwise. Here, $N$ is the total number of policy exception samples.
To enhance the robustness against unbalanced distributions of safe and unsafe data, we integrate PER with balanced accuracy. The combined metric, termed the Policy Exception Score (PES), is defined as the harmonic mean of PER and balanced accuracy and measures the overall performance of the safeguard in adhering to policies while maintaining reliability in safety classification.
$\text{PES}=\frac{2\times\text{PER}\times\text{Acc}}{\text{PER}+\text{Acc}}$ (2)
LlavaGuard Training. We make our training pipeline publicly available. For hyper-parameters and additional details on model tuning, see App. B. We present two different LlavaGuard models with varying model sizes, namely 0.5B and 7B parameters. Both models build upon the Llava-OneVision models (Li et al., 2024b) of their respective sizes.
Inference speed. Speed is a crucial factor when annotating images at scale. The 0.5B model is 347% faster than the 7B model, with an inference time of 0.075s/sample as measured on a single A100 GPU. This speed advantage becomes even more pronounced as GPU memory increases.